OpenStudy (anonymous):

solve for x: 2/5 + 3/5x = x +5/10 Answer: NO SOLUTION

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

the equation is \[\Large \frac{2}{5} + \frac{3}{5x} = \frac{x+5}{10}\] right?

OpenStudy (whpalmer4):

Looking at the solutions, I think it is perhaps more likely to be \[\frac{2}5 + \frac{3}{5}x = \frac{x+5}{10}\]

OpenStudy (anonymous):

WolframAlpha.com came up with\[\frac{2}{5}+\frac{3}{5} x=x+\frac{5}{10} \]where x = -1/4

OpenStudy (whpalmer4):

solution to mine is x = 1/5. @InsanelyChaotic what is the correct form of the equation?

OpenStudy (anonymous):

@whpalmer4 the way @jim_thompson5910 wrote it was correct

jimthompson5910 (jim_thompson5910):

well \[\Large \frac{2}{5} + \frac{3}{5x} = \frac{x+5}{10}\] has two solutions, so the answer isn't "no solutions"

jimthompson5910 (jim_thompson5910):

so double check your work

OpenStudy (anonymous):

-1/4 and -2?

OpenStudy (whpalmer4):

Multiply through by 10x, giving \[4x + 6 = x(x+5)\]\[4x+6=x^2+5x\]\[x^2+x-6=0\]Can you factor that?

OpenStudy (anonymous):

no...

OpenStudy (whpalmer4):

Well, you'll have \[(x-a)(x+b)\]where a*b =6 and the difference between a and b is 1. Can you think of a pair of numbers that makes that true?

OpenStudy (whpalmer4):

What are the factors of 6?

OpenStudy (anonymous):

3

OpenStudy (whpalmer4):

That's one factor, what's another?

OpenStudy (anonymous):

12

OpenStudy (whpalmer4):

Um, no. 12 is not a factor of 6. Factors are the numbers which when multiplied together give you the number you are factoring. What do you have to multiply 3 by to get 6?

OpenStudy (whpalmer4):

___ * 3 = 6 What goes in the blank?

OpenStudy (anonymous):

2

OpenStudy (whpalmer4):

Good. Other factors of 6 are 1 and 6. So, we need two of those factors which when multiplied give 6, and are only 1 apart (difference is 1). Make your selection, please...

OpenStudy (whpalmer4):

Choices are 1, 2, 3, 6

OpenStudy (anonymous):

2

OpenStudy (whpalmer4):

We need two numbers. Not one. The two numbers must multiply to 6, and the difference between them must be 1.

OpenStudy (anonymous):

2 an 3

OpenStudy (whpalmer4):

Very good. So, the middle term is +x, so we need the bigger of the two to be the one with the plus sign (it will be 3x - 2x = 1x): Our equation factors to\[(x+3)(x-2)= 0\]For a product to be 0, one or more of the things being multiplied must be 0. We don't know which, and we want to find all the possible values of x that make that true. So, we set each one equal to 0 and solve for x. \[x+3=0\] \[x-2=0\] What are the two values of x that satisfy those equations? Those are the solutions to this problem.

OpenStudy (whpalmer4):

Checking our factoring: \[(x+3)(x-2) = x(x-2) + 3(x-2) = x^2 -2x + 3x -6 = x^2 + x -6\checkmark\]